A bobsled finishes a race traveling at 32 m/s. After braking the sled slows with an acceleration of -2m/s^2.
How long does it take for the bobsled to stop?
How far does the bobsled travel in the time it takes to stop?
Answer #1

a)
t= v/a
t = 32/2
t = 16 sec .

b) d = vt + 0.5*a*t2
d= 32x16 - 162
d= 256 m .